OpenStudy (anonymous):

1.ac - bc + ad - bd when you group them you add like terms right? but it wants me to factor them in groups how would you do that ?

OpenStudy (phi):

put parens around the first pair and the second pair like this (ac-bc)+(ad-bd) now factor c out of the first pair and factor d out of the second pair. Can you do this?

OpenStudy (anonymous):

what you mean factor like minus out the cs and the ds ?

OpenStudy (phi):

there is a "rule" a(b+c)= (ab+ac) or the other way round: (ab+ac)= a(b+c) Here is the rule using numbers: 2(3+4)= (2*3+2*4) we know it works: 2(3+4)= 2*7=14 and (2*3+2*4)= (6+8)= 14 factoring out "a" means do (ab+ac)= a(b+3c)

OpenStudy (anonymous):

so the anwser would be c(a-)+d(a-b)

OpenStudy (phi):

you mean c(a-b)+d(a-b) right? now the neat thing: factor (a-b) out (pretend is just another letter)

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

so then the factor would be cd?

OpenStudy (phi):

don't ignore the + sign

OpenStudy (anonymous):

oops meant c+d

OpenStudy (phi):

yes. you should get (c+d)(a-b) or the other way round: (a-b)(c+d)